OpenStudy (anonymous):

is there a rational non zero value for b that allows you to factor x^2+b^2? Explain. pretty pretty please help me T^T

OpenStudy (amistre64):

i wonder if we can assume that there is, and work a contradiction from that

OpenStudy (amistre64):

\[(x+\frac nm)(x+\frac nm)=x^2+2\frac nmx+(\frac nm)^2\] \[(x-\frac nm)(x+\frac nm)=x^2+(\frac nm-\frac nm)x-(\frac nm)^2\] \[(x-\frac nm)(x-\frac nm)=x^2-2\frac nmx+(\frac nm)^2\] since none of the possible cases results in \(\large x^2+(\frac nm)^2\) then our assumption is contradicted

OpenStudy (anonymous):

THANKS A BUNCH :D